OpenStudy (anonymous):

ALGEBRA 2 Please help! :) Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. 7. -4, 1,2 8. 2,3,-1

OpenStudy (campbell_st):

well what you need to know is that is x = -4 is a zero then x + 4 is a factor... so find the other 2 factors... then multiply them together for the polynomial. in each question you are dealing with degree 3 polynomials.

OpenStudy (dayakar):

\[\alpha,\beta,\gamma \] are three zeroes of the polynomials then \[f(x) = (x-\alpha)(x-\beta)(x-\gamma)\]

OpenStudy (dayakar):

7)\[\alpha= -4,\beta=1,\gamma=2\] substitute the values and get the polynomial